# Effect of Adaptive Communication Support on Human-AI Collaboration

### v1

@@ Introduction
In this study, we recruited 16 participants (9 male and 7 female) aged between 23 and 33 to evaluate the performance of collaborative agents. We selected participants with varying levels of familiarity with digital agents in game environments: 10 participants with self-reported video game time between 1-10 hours per week, 5 participants reported 10-20 hours per week, and 1 reported 20-30 hours per week. Additionally, 13 of the 16 participants reported prior familiar with Large Language Models (LLMs) and embodied language agents, while 3 of the 16 participants reported no prior knowledge or experiences with language agents.
Procedures
@@
Participants were first asked to complete a consent form per Institutional Review Board (IRB) protocol. Subsequently, prior to the formal trials they were given an opportunity to play with the four agents in an introduction map (Fig. 3) to explore various language feedback styles. During this introduction phase participants were guided through system operations, game rules, and agent functionalities. After the introduction phase, participants proceeded to independently collaborate with each of the four agent types, on easy, medium, and hard layouts (Fig. 3), to complete the maximum score within the given time limit (60 s for each layout). For each participant, the scenario (i.e., agent type $\times$ layout difficulty) was set to show up in a randomized order. We collected a total of 192 experiment trials, 12 trials per participant.
@@
We collected game scores and step-by-step action logs for each experiment shown in (Fig. 3). After completing the teaming scenario, participants were also asked to fill out a survey rating their satisfaction, engagement, and trust level for each experiment, on a seven-point Likert scale. Participants were also asked to specify their preferred language feedback level for each layout.
@@
Additionally, participants were asked to provide improvement suggestions on the language feedback provided by agents (e.g., “If you were to play with this agent in an Overcooked game competition, what changes or improvements would you suggest for the agent’s feedback?”), and state the reason for their satisfaction ratings. For more details, the full questionnaire is available at  [^1] Link to the questionnaire:
https://docs.google.com/forms/d/1xSRPH1mkrZxuKXQ0xvjl9Cv4yi0l-aBblpwyG5MbX34/viewform?edit_requested=true
Results and Discussion
@@
The purpose of our data collection and analysis was to test the following hypotheses: An increased level of language support will result in an increase in the perceived level of trustworthiness and intelligence of the agent, and improve overall team performance.
@@
Surprisingly, our data suggested while language feedback could indeed facilitate human trust, perceived agent intelligence, and team efficiency, the desired level of language support exhibited a different relationship than hypothesized. We report our findings in the subsequent sections.
Language feedback builds human trust and perceived intelligence
@@
Trust is a key factor in human-AI collaboration, shaping human experiences and significantly influencing long-term collaboration efficiency (?). In this section, we report the perceived levels of trust and intelligence after participants teamed with four types of agents across three layouts. Both intelligence and trust levels were measured using a 7-point Likert scale, ranging from “Very Untrustworthy” (or “Very Unintelligent”) to “Very Trustworthy” (or “Very Intelligent”). We found that as the agent’s support level increased from Inactive to Superactive, intelligence rating increased monotonically (Fig. 4). A similar trend was observed in the trust ratings. This suggested that as hypothesized, active communication can facilitate building trust and preceived intelligence. Another interesting observation was that, as the agents become more active, the standard deviation for trust ratings and intelligent ratings also increased. This suggested that both trust and intelligence levels also become more influenced by individual human preferences when the agent takes on a more active role in assisting.
@@
Figure 4: Human perceived agent intelligence level (blue bar) and trust level (red bar) represented on a seven-point Likert scale, ranging from “Very Unintelligent/Untrustworthy” to “Very Intelligent/Trustworthy”.
Appropriate Language feedback improves collaboration efficiency
@@
We used the game score to evaluate the team perfromance and collaboration efficiency between humans and agents. Overall, the team scored more points in easy layout with all agent types (53.6 in average) as compared to the medium (30.6 in average) and hard (27.3 in average) layouts. This is not surprising, as the easy layout has simpler maps and fewer subtasks, requiring minimal coordination to complete. In contrast, medium and hard layouts introduced more complex subtasks and dependencies, requiring greater coordination and team effort, which could increase the chance of errors during the task and result in lower scores.
@@
Figure 5: Game scores of all participants paired with different agents across various layouts. The score distribution of each agent type, IFA, PFA, AFA, SFA, is represented by the blue, red, green, and yellow boxes, respectively.
However, the team performance was not always better with more active agents. In the easy layout, team with PFA performed best, scoring 83.6 in average, significantly higher than IFA, AFA, and SFA (Fig. 5). It was expected that PFA performed better than IFA in the easy task, as it provided effective support with minimal interference, enabling participants to benefit from its assistance as needed and allowing humans to take the lead. This was also supported by user feedback from the survey – “Robot listening to the user and following the commands would help the game better. In terms of intelligence, passive robots work better but still lack user assistance”. Unexpectedly, while AFA and SFA offered more active feedback, they achieved lower scores than the PFA (Fig. 5), suggesting that constant communication was less effective for simple tasks, and may even distract humans. As the complexity increases to the medium level, the team with SFA exhibited a huge increase in score (Fig. 5), exceeding the performance of PFA, implying that as task complexity increased, the frequent support provided by SFA became more valuable. Similarly, in the hard layout, the SFA and AFA demonstrated superior performance as compared to IFA and PFA (Fig. 5), underscoring the value of active guidance and high-frequency support in facilitating collaboration and helping improve task execution in complex tasks.
Interestingly, despite SFA’s higher frequency of active support compared to AFA, overall performance still decreased due to the high complexity of the hard layout compared to AFA. Participant feedback highlighted this challenge, e.g.,the robot/agent is not as smart as me,” and, “I have to give it a long instruction set, and I am more intelligent than it in the hard layout.” One interpretation of these responses was that participants found the agent’s support insufficient for complex tasks. As a result, instead of reducing cognitive load, the frequent suggestions from the agent required human to carefully think about responses, which ultimately increased cognitive demands and decreased team efficiency. Another interpretation is that psychologically, humans may have the preference to demonstrate and maintain intellectual superiority in challenging tasks when teaming with AI agents. As a result, how AI agents communicate suggestions may greatly influence humans’ acceptance rate and team efficiency.
@@
Humans don’t always prefer the best-performing agents
@@
We believe this preference shift can be explained by the flow theory (??), which states that people feel most engaged when task complexity aligns with their skill level, and robots can help achieve this balance by adjusting their level of support. In the easy layout, the task was not significantly beyond human’s skill level, and the need for additional help to maintain engagement was low. Therefore, while the active agents can provide help, this help did not significantly influence human’s engage level and satisfaction rate (Fig. 6B). As a result, the non-active agent, which requires the lowest level of cognitive load (“cost”) was preferred. As the difficulty level increases, however, the gap between task complexity and human skill level increases, resulting a disruption to the task-capability balance and reduced engagement. Meanwhile, the additional feedback and support from the more active agents, such as assigning subtasks (e.g., “pick up the onion from (x, x).”, can reduce the coordination and planning effort on the human, and restore the task-capability balance and enhance human’s feel of engagement and joy. At this point, the cost of communication became negligible as compared to the need for sense of achievement, making more active agents more desirable in these challenging scenarios (Fig. 6B).
@@
To test this, we recorded participants’ engagement during each game trial. As shown in Fig. 7, satisfaction levels increased with higher engagement, with a correlation test yielding a p-value of 0.00275 and a correlation coefficient of 0.93, indicating a strong relationship. This strong correlation between satisfaction and engagement indicates that engagement is the cause of different satisfaction levels reported by participants. This aligned with our theory, emphasizing the need to align agent support and feedback frequency based on task complexity relative to human capabilities.
@@
Adaptively assigning subtasks and providing language feedback
Our findings revealed the importance for LLM agents to adapt their language feedback by considering the relationship between task complexity, $T_{h}$, human capability, $C_{h}$, and the LLM’s capability, $C_{l}$. Below we propose a simple adaptation strategy for LLM agents to select their support level and language feedback frequency:
  • $C_{h}>T$ and $C_{l}<T$: Here human capability surpasses the task’s complexity, and the LLM capability is not sufficient to address the challenging tasks without human guidance. Based on our results, a passive (PFA) to relative infrequent (AFA) feedback style would allow the agent to provide sufficient support to improve team performance and request human help when needed, while keeping communication frequency to a minimal to avoid overhead on the human side. This way, human teammates with higher capability level could guide the agent on high-level coordination strategies and specific subtask execution, keeping them engaged.
  • $C_{h}<T$ and $C_{l}>T$: Here task complexity exceeds human capability, while LLM is fully capable of executing the task. In this case, extra active agent feedback and support (SFA) are crucial for maintaining team performance, and help reduce the gap between human capability and task complexity.
  • $C_{h}<T$ and $C_{l}<T$: Here task complexity exceeds human capability. However, LLM capability is also not sufficient to address the challenging task neither. In this case, language feedback from LLM is often not useful in resolving the challenges that human is struggling with, and can even be misleading. High-frequency communications from LLM in this scenario would require additional effort from human to respond, potentially further heightening the anxiety that human is already experiencing (?), decreasing team performance and human engagement. Our results suggested that a more passive (PFA) or relative infrequent (AFA) feedback style would result in better teaming performance in this case.
  • $C_{h}>T$ and $C_{l}>T$: Here both human and LLM capabilities surpass the task’s complexity. The active feedback style (AFA) could allow the human and agents to communicate their needs at a comfortable pace, and improve collaboration efficiency.
Conclusion
@@
Our user study results demonstrated that language-based feedback from LLMs can significantly enhance collaboration performance and foster human trust, and that as task complexity increases, more frequent, proactive support is desired. However, our study also revealed that it is critical for the agent to select their language feedback frequency based on task complexity, human capability, and agent capability. Overly frequent feedback in simple tasks or from less capable agents does not effectively increase team performance and satisfaction, and could even increase effort and reduce human engagement. Based on these findings, we proposed a simple principle that allows agents to adapt their language feedback style according to perceived task challenges, human capabilities, and LLM capabilities.
Limitations and future work
In this work, we focused on investigating the effect of agent feedback frequency on team performance. As such, each agent was set to a constant active level, and cannot dynamically adjust their level of support and language feedback throughout the task. In real-world scenarios, task complexity and human skill levels for different sub task can vary dynamically, requiring agents to adjust their communications and behaviors accordingly. Future work should explore methods for adaptive agent to estimate human cognitive load, capabilities, and engagement, and adjust LLM feedback in real time to enable better team performance and adaptability. The results from our study provide the basis for designing and implementing such real-time feedback adjustments. Going forward, these adaptive response and feedback capabilities can empower future LLM agents to flexibly support human needs in a wide variety of task complexities, fostering true partnerships and enhancing teamwork outcomes.
@@ Acknowledgment
This work is supported by the NASA Planetary Science and Technology Through Analog Research (PSTAR) program, Award # 80NSSC22K1313. The authors would like to thank Stefanos Nikolaidis for helpful discussions, and Siddharth Srikanth for assistance with prompt engineering and code refactoring.

### v2

@@ 5.2 Participants
In this study, we recruited 16 participants (9 male and 7 female) aged between 23 and 33 to evaluate the performance of collaborative robotic agents. We selected participants with varying levels of familiarity with digital agents in-game environments: 10 participants self-reported video game time between 1-10 hours per week, 5 participants reported 10-20 hours per week, and 1 reported 20-30 hours per week. Additionally, 13 of the 16 participants reported prior familiarity with LLMs and embodied language agents, while 3 of the 16 participants reported no prior knowledge or experiences with language agents.
5.3 Procedures
@@
Participants were first asked to complete a consent form per the Institutional Review Board (IRB) protocol. Subsequently, prior to the formal trials they were given an opportunity to play with the four robotic agents in an introduction map (Fig. 3) to explore various language feedback styles. During this introduction phase participants were guided through system operations, game rules, and robot functionalities. After the introduction phase, participants proceeded to independently collaborate with each of the four agent types on easy, medium, and hard layouts (Fig. 3) to complete the maximum score within the given time limit (60 s for each layout). For each participant, the scenario (i.e., agent type $\times$ layout difficulty) was set to show up in a randomized order. We collected a total of 192 experiment trials, 12 trials per participant.
@@
We collected game scores for each experiment shown in (Fig. 3). After completing the teaming scenario, participants were also asked to fill out a survey rating their satisfaction, engagement, and trust level for each experiment, on a seven-point Likert scale. Participants were also asked to specify their preferred language feedback level for each layout.
@@
Additionally, participants were asked to provide improvement suggestions on the language feedback provided by robots (e.g., “If you were to play with this robotic agent in an Overcooked game competition, what changes or improvements would you suggest for the agent’s feedback?”), and state the reason for their satisfaction ratings. For more details, the full questionnaire is available in the Appendix. 11.
6 Results and Discussion
@@
The purpose of our data collection and analysis was to test the following hypotheses: An increased level of language support will result in an increase in the perceived level of trustworthiness and intelligence of the robot, and improve overall team performance.
@@
Surprisingly, our data suggested while language feedback could indeed facilitate human trust, perceived robot intelligence, and team efficiency, the desired level of language support exhibited a different relationship than hypothesized. We report our findings in the subsequent sections.
6.1 Language Feedback Builds Human Trust and Perceived Intelligence
@@
Trust is a key factor in human-AI collaboration, shaping human experiences and significantly influencing long-term collaboration efficiency [8]. In this section, we report the perceived levels of trust and intelligence after participants teamed with four types of robotic agents across three layouts. Both intelligence and trust levels were measured using a 7-point Likert scale, ranging from “Very Untrustworthy” (or “Very Unintelligent”) to “Very Trustworthy” (or “Very Intelligent”). We found that as the robot’s support level increased from Inactive to Superactive, intelligence rating increased monotonically (Fig. 4). A similar trend was observed in the trust ratings. This suggested that as hypothesized, active communication can facilitate building trust and preceived intelligence. Another interesting observation was that, as the robot become more active, the standard deviation for trust ratings and intelligent ratings also increased. This suggested that both trust and intelligence levels also become more influenced by individual human preferences when the robot takes on a more active role in assisting.
@@
Figure 4: Human perceived robot intelligence level (blue bar) and trust level (red bar) represented on a seven-point Likert scale, ranging from “Very Unintelligent/Untrustworthy” to “Very Intelligent/Trustworthy”.
6.2 Appropriate Language Feedback Improves Collaboration Efficiency
@@
We used the game score to evaluate the team performance and collaboration efficiency between humans and robots. Overall, the team scored more points in the easy layout with all agent types (53.6 on average) as compared to the medium (30.6 on average) and hard (27.3 on average) layouts. This is not surprising, as the easy layout has simpler maps and fewer subtasks, requiring minimal coordination to complete. In contrast, medium and hard layouts introduced more complex subtasks and dependencies, requiring greater coordination and team effort, which could increase the chance of errors during the task and result in lower scores.
@@
Figure 5: Game scores of all participants paired with different robotic agents across various layouts. The score distribution of each agent type, IFA, PFA, AFA, SFA, is represented by the blue, red, green, and yellow boxes, respectively.
However, the team performance was not always better with more active robotic agents. In the easy layout, team with PFA performed best, scoring 83.6 in average, significantly higher than IFA, AFA, and SFA (Fig. 5). It was expected that PFA performed better than IFA in the easy task, as it provided effective support with minimal interference, enabling participants to benefit from its assistance as needed and allowing humans to take the lead. This was also supported by user feedback from the survey – “Robot listening to the user and following the commands would help the game better. In terms of intelligence, passive robots work better but still lack user assistance”. Unexpectedly, while AFA and SFA offered more active feedback, they achieved lower scores than the PFA (Fig. 5), suggesting that constant communication was less effective for simple tasks, and may even distract humans. As the complexity increases to the medium level, the team with SFA exhibited a huge increase in score (Fig. 5), exceeding the performance of PFA, implying that as task complexity increased, the frequent support provided by SFA became more valuable. Similarly, in the hard layout, the SFA and AFA demonstrated superior performance as compared to IFA and PFA (Fig. 5), underscoring the value of active guidance and high-frequency support in facilitating collaboration and helping improve task execution in complex tasks.
Interestingly, despite SFA’s higher frequency of active support compared to AFA, overall performance still decreased due to the high complexity of the hard layout compared to AFA. Participant feedback highlighted this challenge, e.g.,the robot/agent is not as smart as me,” and, “I have to give it a long instruction set, and I am more intelligent than it in the hard layout.” One interpretation of these responses was that participants found the robot’s support insufficient for complex tasks. As a result, instead of reducing cognitive load, the frequent suggestions from the robot required human to carefully think about responses, which ultimately increased cognitive demands and decreased team efficiency. Another interpretation is that psychologically, humans may have the preference to demonstrate and maintain intellectual superiority in challenging tasks when teaming with robots. As a result, how robotic agents communicate suggestions may greatly influence humans’ acceptance rate and team efficiency.
@@
6.3 Humans Don’t Always Prefer the Best-performing Robotic Agents
@@
We believe this preference shift can be explained by the flow theory [11, 7], which states that people feel most engaged when task complexity aligns with their skill level, and robots can help achieve this balance by adjusting their level of support. In the easy layout, the task was not significantly beyond human’s skill level, and the need for additional help to maintain engagement was low. Therefore, while the active robotic agents can provide help, this help did not significantly influence human’s engage level and satisfaction rate (Fig. 6B). As a result, the non-active agent, which requires the lowest level of cognitive load (“cost”) was preferred. As the difficulty level increases, however, the gap between task complexity and human skill level increases, resulting a disruption to the task-capability balance and reduced engagement. Meanwhile, the additional feedback and support from the more active robotic agents, such as assigning subtasks (e.g., “pick up the onion from (x, x).”, can reduce the coordination and planning effort on the human, and restore the task-capability balance and enhance human’s feel of engagement and joy. At this point, the cost of communication became negligible as compared to the need for sense of achievement, making more active robotic agents more desirable in these challenging scenarios (Fig. 6B).
@@
To test this, we recorded participants’ engagement during each game trial. As shown in Fig. 7, satisfaction levels increased with higher engagement, with a correlation test yielding a p-value of 0.00275 and a correlation coefficient of 0.93, indicating a strong relationship. This strong correlation between satisfaction and engagement indicates that engagement is the cause of different satisfaction levels reported by participants. This aligned with our theory, emphasizing the need to align robot support and feedback frequency based on task complexity relative to human capabilities.
@@
7 Adaptively assigning subtasks and providing language feedback
Our findings revealed the importance of LLM-powered robots adapting their language feedback by considering the relationship between task complexity, $T_{h}$, human capability, $C_{h}$, and the LLM’s capability, $C_{l}$. Below we propose a simple adaptation strategy for LLM-powered robots to select their support level and language feedback frequency:
  • $C_{h}>T$ and $C_{l}<T$: Here human capability surpasses the task’s complexity, and the LLM capability is not sufficient to address the challenging tasks without human guidance. Based on our results, a passive (PFA) to relative infrequent (AFA) feedback style would allow the robot to provide sufficient support to improve team performance and request human help when needed, while keeping communication frequency to a minimal to avoid overhead on the human side. This way, human teammates with higher capability level could guide the robot on high-level coordination strategies and specific subtask execution, keeping them engaged.
  • $C_{h}<T$ and $C_{l}>T$: Here task complexity exceeds human capability, while LLM is fully capable of executing the task. In this case, extra active robot feedback and support (SFA) are crucial for maintaining team performance, and help reduce the gap between human capability and task complexity.
  • $C_{h}<T$ and $C_{l}<T$: Here task complexity exceeds human capability. However, LLM capability is also not sufficient to address the challenging task neither. In this case, language feedback from LLM is often not useful in resolving the challenges that human is struggling with, and can even be misleading. High-frequency communications from LLM in this scenario would require additional effort from human to respond, potentially further heightening the anxiety that human is already experiencing [21], decreasing team performance and human engagement. Our results suggested that a more passive (PFA) or relative infrequent (AFA) feedback style would result in better teaming performance in this case.
  • $C_{h}>T$ and $C_{l}>T$: Here both human and LLM capabilities surpass the task’s complexity. The active feedback style (AFA) could allow the human and robots to communicate their needs at a comfortable pace, and improve collaboration efficiency.
8 Conclusion
@@
Our user study results demonstrated that language-based feedback from LLMs can significantly enhance collaboration performance and foster human trust, and that as task complexity increases, more frequent, proactive support is desired. However, our study also revealed that it is critical for the robot to select their language feedback frequency based on task complexity, human capability, and robot capability. Overly frequent feedback in simple tasks or from less capable robots does not effectively increase team performance and satisfaction, and could even increase effort and reduce human engagement. Based on these findings, we proposed a simple principle that allows robots to adapt their language feedback style according to perceived task challenges, human capabilities, and LLM capabilities.
9 Limitations
In this work, we focused on investigating the effect of robot feedback frequency on team performance. As such, robots were set to a constant active level and could not dynamically adjust their level of support and language feedback throughout the task. In real-world scenarios, task complexity and human skill levels for different sub-tasks can vary dynamically, requiring robots to adjust their communications and behaviors accordingly. Future work should explore methods for the adaptive robots to estimate human cognitive load, capabilities, and engagement and adjust LLM feedback in real-time to enable better team performance and adaptability. The results from our study provide the basis for designing and implementing such real-time feedback adjustments. Going forward, these adaptive response and feedback capabilities can empower future LLM-powered robots to flexibly support human needs in a wide variety of task complexities, fostering true partnerships and enhancing teamwork outcomes.
@@ Acknowledgment
We omitted the acknowledgment to comply with the double-blind review policy.